Solve by factoring: 2x^2+126=32x
Answer by Alan3354(69443)

Solve by factoring: 2x^2+126=32x
------------
2x^2 - 32x + 126 = 0
x^2 - 16x + 63 = 0
----------
Find a pair of factors of 63 that add to -16 (both negative).